OpenStudy (anonymous):

Two cubes are tossed. Find P(only one shows a 2).

OpenStudy (anonymous):

I'm assuming you're talking about normal dice.

OpenStudy (ddcamp):

P(A die has a 2) AND P(A die doesn't have a 2) If you know each probability, you can find the overall probability.

OpenStudy (anonymous):

You must evaluate the odds that one shows a two and the other doesn't, but the easiest way for this one is just to think about how many options fit the conditions.

OpenStudy (ddcamp):

Remember, the probability of event A and B happening is the product of P(A) and P(B).

OpenStudy (anonymous):

i think it is a binomial probability where probability os suscess = 1/6 probability of failure = 5/6 so the answer is C(2,1)*(1/6)*(5/6)

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Thank you :) Final answer is 5/18

OpenStudy (anonymous):

Also correct!

OpenStudy (anonymous):

I wish I could give you both medals.